# Context parroting: A simple but tough-to-beat baseline for foundation models in scientific machine learning

### v1

@@ 5.1 Context parroting versus foundation models
Figure 2: Context parroting outperforms foundation models in zero-shot forecasting. Left: Forecast error of each model as a function of the forecast horizon. The context length is set to $512$ for all models. Right: Valid prediction time of each model as a function of the context length. Dashed lines indicate context lengths exceeding the maximum context window of the model. All results are averaged over the $135$ chaotic systems in the dysts database, with $20$ trajectories from random initial conditions for each system.
Here, we compare context parroting and foundation models in their ability to predict chaotic dynamics. Figure 2 shows each model’s forecasting error and valid prediction time. It is clear that context parroting is better than all foundation models in both metrics. In Fig. 9, we show that this remains true when benchmarked against MSE and MAE. Among foundation models, Chronos is the best performer in predicting chaotic systems, which is not surprising given that it utilizes parroting as a main forecasting strategy. Chronos’s tendency to context parrot arises from its distinct architecture as a language model that implicitly quantizes time series. As a result, Chronos is trained using cross-entropy loss, which incentivizes preservation of k-gram frequencies, ensuring repetition of sequences in the output and incentivizing generating diverse forecast samples consistent with the dynamical system’s underlying measure. In contrast, TimeMoE and TimesFM are trained using mean squared error loss, and so these models lose diversity and forecast the mean at long forecast horizons. Some representative forecasts from the foundation models are shown in Fig. 8.
@@
Moreover, as we demonstrate in Fig. 3, the inference cost of context parroting is negligible compared to all foundation models (not to mention the substantial GPU time needed to pre-train foundation models). For example, there is an over six orders of magnitude gap between Chronos and context parroting for all context lengths. Combined with the fact that the performance of parroting is not sensitive to the choice of the embedding dimension $D$ (Fig. 4), these results establish context parroting as a simple but hard-to-beat baseline for zero-shot forecasting of dynamical systems.
Figure 4: Effect of the embedding dimension $D$ on the forecast accuracy of context parroting. Overall, the valid prediction time stays consistent over a wide range of embedding dimension $D$. For short context windows, there is a slight advantage to small $D$. For long context windows, larger embedding dimensions are marginally better. This observation suggests potential improvements in the future that choose $D$ adaptively based on factors such as context length.
Figure 5 further explores the effect of context length on forecast accuracy. We find that longer context windows generally lead to better performance for both context parroting and Chronos. However, the longest context length Chronos can effectively utilize is $512$ data points; anything beyond that will only increase inference cost without improving accuracy. This limit is determined by Chronos’s maximum context window chosen at pre-training. To be able to utilize longer context, Chronos must be retrained from scratch with much more data and compute. Context parroting, on the other hand, is happy to utilize context data of any length. Interestingly, Chronos outperforms context parroting on short contexts, which points to additional zero-shot learning strategies beyond parroting employed by Chronos. This is perhaps not surprising given that at short context length, the times series becomes effectively nonstationary, which is the strength of time-series foundation models. Moreover, even when restricted to parroting, Chronos can in principle dynamically choose the optimal embedding dimension $D$ for each individual time series, giving it an advantage over parroting algorithms with a fixed $D$. It would be interesting to explicitly identify the mechanisms that enable Chronos to outperform parroting in this regime.
@@
Figure 5: Parroting can better utilize longer context data while Chronos does better for shorter contexts. Each line represents a different context length. The performance of Chronos saturates once the context length exceeds $512$ data points, whereas the accuracy of context parroting keeps improving for longer context windows. Here we set the embedding dimension $D=10$ for the parroting algorithm.
@@ 5.2 In-context neural scaling law
Liu et al. [61] reported an in-context neural scaling law for LLMs applied to learning dynamical systems, in which the one-step forecast error decreases algebraically with context length. However, it is unclear where this scaling law came from or why LLMs trained on text can be effective for time series without fine tuning. Here, we show that context parroting naturally gives rise to the same in-context scaling law and provides geometric insights into its origin. Given the similarity between parroting and the induction heads implemented by LLMs [62], the geometric explanation we develop next for context parroting can conceivably be applied to LLMs and explain the observations in Liu et al. [61].
The left panel in Fig. 6 reproduces the power law relation between one-step forecast error (measured by sMAPE) and context length for the parroting method. Why do longer context lengths enable better predictions of chaotic systems? This is because more context data allows the algorithm to find better matching motifs, and a closely matched motif allows the parroted sequence to shadow the ground truth for longer. The overlap between the matching motifs can be measured by their Euclidean distance. For length-$D$ motifs, this is equivalent to embed the context trajectory in a $D$-dimensional delay-embedded space (mapping $x_{s}$ to $x_{s-D+1:s}$) and find the distance between the embedded last context point $x_{t-D+1:t}$ and its nearest neighbor. The right panel in Fig. 6 shows the improving overlap explicitly, with the distance between the matching motifs decreasing algebraically with context length.
Figure 6: Scaling laws with context length. Left: One-step forecast error vs context length. The scaling follows a power law for all embedding dimensions considered. Smaller $D$ is more accurate here because the use of the one-step forecast error. Larger $D$ can be more accurate for longer forecasting horizons. Right: Euclidean distance between the last context motif $x_{t-D+1:t}$ and its nearest neighbor in the $D$-dimensional delay-embedded space as a function of context length. Again, the scaling follows a power law for all $D$. The forecast accuracy is directly tied to the embedding distance: Smaller distances translate into better predictions. As the number of data points is increased, it becomes more and more likely that a context motif will land in the vicinity of the last context motif, and the correlation dimension of the chaotic attractor determines the rate of approach. In principle, infinite context length should give context parroting infinite accuracy for any stationary and deterministic systems.
@@
where $C(\epsilon)$ is the number of point pairs in the attractor that are within a given radius $\epsilon$ from each other. In other words, if we plot $C(\epsilon)$ against $\epsilon$ on a log-log plot, $d_{\mathrm{cor}}$ would be the slope of the plotted line. Due to the ergodic property of chaotic attractors, the context trajectory can be seen as a random sample of the attractor. Longer context trajectory contains more samples, and the expected distance between two context points in a delay-embedded space decreases with context length as $L^{-1/d_{\mathrm{cor}}}$. For example, for a two-dimensional attractor, the distance between two random points on the attractor will decrease as $1/\sqrt{L}$. Correlation dimension thus measures the speed at which the minimum embedding distance between points on an attractor can be reduced by including more samples, and higher dimensions require more points to reduce the distance to the same extent. Mathematically, we thus expect $\alpha=1/d_{\mathrm{cor}}$. However, numerically it is challenging to accurately estimate the scaling coefficient $\alpha$ due to noise in the data. Nonetheless, in Fig. 7, we observe strong correlation between $d_{\mathrm{cor}}$ and $1/\alpha$, supporting our theoretical argument above.
Figure 7: Strong correlation between the inverse scaling coefficient $1/\alpha$ and the correlation dimension $d_{\mathrm{cor}}$ of the chaotic attractor. We estimate $\alpha$ from the $\ell$ vs $L$ plot obtained with context parroting at $D=10$, where $L$ ranges from $2^{6}$ to $2^{16}$. This is done separately for each individual systems. Thus, each dot represents one of the $135$ chaotic systems included in our benchmark. The Spearman rank-order correlation coefficient is $0.44\pm 0.08$ (bootstrapped standard error), and a linear regression with standard error range is underset (shaded interval). The dashed line corresponds to an exact $1:1$ scaling.
@@ 6 Conclusion and future directions
We find that a simple forecast strategy outperforms leading foundation models on dynamical systems forecasting, a critical task in scientific machine learning. This surprising finding exposes a limitation of current time-series foundation models as general-purpose time-series forecasters and highlights the need to further scale them or to fine-tune them for specific domains. It also suggests that accurately measuring the performance of foundation models can be difficult for scientific machine learning tasks, because simple, uninformative strategies like parroting can effectively game both short- and long-term accuracy metrics.
Finding a simple but tough-to-beat baseline for a challenging task can encourage rethinking of the status quo, motivating better model architectures and baselines [80]. For example, identifying zero-shot learning strategies beyond parroting can spur the development of next-generation foundation models and contribute to the debate on whether (or to what extent) large language models are stochastic parrots [81, 82, 83, 84]. Context parroting formalizes an explicit baseline to compare against in the time-series domain and can help discover beyond-parroting strategies.
An important future direction is to generalize context parroting to deal with non-stationary time series while keeping the simplicity and efficiency of the method. A promising avenue involves combining parroting with a probabilistic model, such as a Gaussian Process, to account for nonstationary trends. Additionally, the diversity of long-term forecasts can be improved by allowing stochastic selection among multiple candidate forecasts. Once generalized, the non-stationary parroting method can replace Naive and Seasonal Naive to serve as a more informative baseline for the zero-shot forecasting of general time series (weather, traffic, finance, etc.).
@@ Appendix A Sample predictions from foundation models
Figure 8: Example forecasts on a chaotic system from foundation models. This is the same task as presented in Fig. 1 (predicting the $x$ variable of the Lorenz system based on a short context trajectory with $512$ data points). Chronos does extremely well with a parroting strategy. The other models perform comparatively poorly and all exhibit a tendency to underestimate the oscillations (e.g., by quickly converging towards the mean). This is a general trend that we consistently observe across different chaotic systems and initial conditions.
@@ Appendix B Benchmarking with other metrics
Figure 9: Context parroting outperforms foundation models in zero-shot forecasting. Same setup as in Fig. 2, but with the forecast error measured by MSE (left) and MAE (right).
Appendix C Context parroting
C.1 Discrete-token parroting

### v2

@@ 5.1 Context parroting versus foundation models
Figure 2: Context parroting outperforms foundation models in zero-shot forecasting for both short-term point-wise accuracy and long-term attractor reconstruction. Left: Forecast error of each model as a function of the forecast horizon. The context length is set to $512$ for all models. Right: KL Divergence between the predicted attractors and the true attractors (smaller is better). Solid lines represent mean and dashed lines represent median. All results are obtained from $135$ chaotic systems in the dysts database, with 2020 trajectories from random initial conditions for each system.
Here, we compare context parroting and foundation models in their ability to predict chaotic dynamics. Figure˜2 shows each model’s forecasting error (measured by sMAPE) as well as their accuracy in attractor reconstruction (measured by KL Divergence). It is clear that context parroting is better than all foundation models tested here in both metrics. In Fig.˜7, we show that this remains true when benchmarked against MSE and MAE. The results for correlation dimension accuracy are shown in Fig.˜8, where parroting and Chronos significantly outperform the rest of the foundation models. We also explore the effects of observational noise and data granularity on the forecast performance in Tables˜4, 6, 5, 7, 8, 9, 11, 10, 13 and 12. Over the wide range of noise intensity and data granularity we tested, parroting is always the best or the second best according to all metrics (VPT, MAE, MSE, correlation dimension accuracy, and attractor KL Divergence).
Among foundation models, Chronos is the best performer in predicting chaotic systems, which is not surprising given that it utilizes parroting as a main forecasting strategy [42]. Chronos’s tendency to context parrot arises from its distinct architecture as a language model that implicitly quantizes time series. As a result, Chronos is trained using cross-entropy loss, which incentivizes preservation of k-gram frequencies and encourages the generation of diverse forecast samples consistent with the dynamical system’s underlying measure. In contrast, TimeMoE and TimesFM are trained using mean squared error loss. As a result, these models lose diversity and forecast the mean at long forecast horizons (i.e., they tend to underestimate the oscillations). Some representative forecasts from the foundation models are shown in Fig.˜6, which shows “regressing-to-the-mean” as a common failure mode for many foundation models when forecasting chaotic systems.
@@
Moreover, as we demonstrate in Fig.˜3, the inference cost of context parroting is negligible compared to all foundation models (not to mention the substantial GPU time needed to pre-train foundation models). For example, there is an over six orders of magnitude computational gap between Chronos and context parroting for all context lengths. Combined with the fact that the performance of parroting is not sensitive to the choice of the embedding dimension $D$ (Fig.˜10), these results establish context parroting as a simple but effective baseline for zero-shot forecasting of dynamical systems.
@@
Figure 4: Parroting can better utilize longer context data while Chronos does better for shorter contexts. Each line represents the forecast error for a different context length. The performance of Chronos saturates once the context length exceeds its designed upper limit of $512$ data points, whereas the accuracy of context parroting keeps improving for longer context windows. Here we set the embedding dimension $D=10$ for the parroting algorithm.
Figure˜4 further explores the effect of context length on forecast accuracy. We find that longer context windows generally lead to better performance for both context parroting and Chronos. However, the longest context length Chronos can effectively utilize is $512$ data points; anything beyond that will only increase inference cost without improving accuracy. This limit is determined by Chronos’s maximum context window chosen at pre-training. To be able to utilize longer context, Chronos must be retrained from scratch with much more data and compute. Context parroting, on the other hand, is happy to utilize context data of any length. In Fig.˜9, we show VPT as a function of the context length for context parroting and all four foundation models. Again, parroting is the clear winner for medium to long context lengths, easily reaching an average VPT of over $5$ Lyapunov time when given long enough context.
Interestingly, Chronos outperforms context parroting on short contexts, which points to additional zero-shot learning strategies beyond parroting employed by Chronos. This is perhaps not surprising given that at short context length, the time series becomes effectively nonstationary, which is the strength of time-series foundation models. For example, Chronos is great at continuing the local trend in the context, which can be a more effective strategy than parroting when the length of the context is limited. Moreover, even when restricted to parroting, Chronos can in principle dynamically choose the optimal embedding dimension $D$ for each individual time series, giving it an advantage over parroting algorithms with a fixed $D$. It would be interesting to explicitly identify the mechanisms that enable Chronos to outperform parroting in the short-context regime.
@@ 5.2 In-context neural scaling law
Liu et al. [62] reported an in-context neural scaling law for LLMs applied to dynamical systems, in which the one-step forecast error decreases algebraically with context length. However, it is unclear where this scaling law came from or why LLMs trained on text can be effective for time series without fine tuning. Here, we show that context parroting naturally gives rise to the same in-context scaling law and provides geometric insights into its origin. Given the similarity between parroting and the induction heads implemented by LLMs [64], the geometric explanation we develop next for context parroting can conceivably be applied to LLMs and explain the observations in Liu et al. [62].
Figure 5: Scaling laws with context length. Left: One-step forecast error versus context length. The scaling follows a power law for all embedding dimensions $D$ considered. Smaller $D$ is more accurate here because the use of the one-step forecast error. Larger $D$ can be more accurate for longer forecasting horizons. Right: Euclidean distance between the last context motif $x_{L-D+1:L}$ and its closest match, as a function of context length. Again, the scaling follows a power law for all $D$. The forecast accuracy is directly tied to the motif distance: Smaller distances translate into better predictions. As the number of context data points is increased, it becomes more and more likely that a context motif will land in the vicinity of the last context motif, and the correlation dimension of the chaotic attractor determines the rate of approach. In principle, infinite context length should give context parroting infinite accuracy for any deterministic system with a well-defined attractor.
The left panel in Fig.˜5 shows the power law relation between one-step forecast error (measured by sMAPE) and context length for the parroting method. Why do longer context lengths enable better predictions of chaotic systems? This is because more context data allows the algorithm to find better matching motifs, and a closely matched motif allows the parroted sequence to shadow the ground truth for longer. The overlap between the matching motifs can be measured by their Euclidean distance. For length-$D$ motifs, this is equivalent to embed the context trajectory in a $D$-dimensional delay-embedded space (i.e., mapping $x_{s}$ to $x_{s-D+1:s}$) and find the distance between the embedded last context point $x_{L-D+1:L}$ and its nearest embedded neighbor. The right panel in Fig.˜5 shows the improving overlap explicitly, with the distance between the matching motifs decreasing algebraically with context length. It is easy to see that embedding distance should map linearly to the expected one-step forecast error, so the power law for the embedding distance implies the power law for one-step forecast error, as confirmed by Fig.˜5.
@@
where $C(\epsilon)$ is the number of point pairs in the attractor that are within a given radius $\epsilon$ from each other. In other words, if we plot $C(\epsilon)$ against $\epsilon$ on a log-log plot, $d_{\mathrm{cor}}$ would be the slope of the plotted line. Due to the ergodic property of chaotic attractors, the context trajectory can be seen as a random sample of the attractor. Longer context trajectory contains more samples, and the expected distance between two context points in a delay-embedded space decreases with context length as $L^{-1/d_{\mathrm{cor}}}$. For example, for a two-dimensional attractor, the distance between two random points on the attractor will decrease as $1/\sqrt{L}$. Correlation dimension thus measures the speed at which the minimum embedding distance between points on an attractor can be reduced by including more samples, and higher dimensionality requires more points to reduce the distance to the same extent. Mathematically, we thus expect $\alpha=1/d_{\mathrm{cor}}$. A similar scaling law has been derived for the Farmer-Sidorowich forecasting method from the nonlinear dynamics community [77]. Despite the theoretical correspondence, however, numerically it is challenging to accurately estimate the scaling coefficient $\alpha$ due to noise in the data. Nonetheless, in Fig.˜11, we observe strong correlation between $d_{\mathrm{cor}}$ and $1/\alpha$, supporting our theoretical argument above.
5.3 SciML tasks beyond low-dimensional chaotic systems
So far we focused on low-dimensional chaotic systems from the dysts dataset, which enabled systematic comparison between different forecasting models with standardized benchmarks. Here, we show that parroting also outperforms foundation models on a broader class of SciML tasks, including real-world datasets of current scientific interest. Our datasets cover fields such as fluid dynamics (von Karman vortex street with Reynolds number $\mathrm{Re}=900$, decomposed into PCA modes), neuroscience (ECG recordings from the QT Database in PhysioNet), engineering ($28$ coupled electronic circuits measured experimentally, from [86]), and physics ($23$ Kuramoto oscillators coupled through frustrated and nonreciprocal interactions, recently studied in [87]). These are all high-dimensional systems, two generated from simulations and two measured in the real world. For the metrics, we use MAE and MSE to measure pointwise forecast accuracy, and KL Divergence to measure the accuracy in attractor reconstruction. The results are summarized below. Parroting is either the best or the second best in all experiments. Other metrics, such as valid prediction time and correlation dimension accuracy, give similar rankings.
Task Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 1: Performance comparison (MAE @ 50 steps) of forecasting models across SciML tasks. Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Task Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 2: Performance comparison (MSE @ 50 steps) of forecasting models across SciML tasks. Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Task Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 3: Performance comparison (KL Divergence between predicted and true attractors) of forecasting models across SciML tasks. Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
@@ 6 Conclusion and future directions
We find that a simple forecast strategy—context parroting—outperforms leading foundation models on dynamical systems forecasting, a critical task in scientific machine learning. This surprising finding exposes a limitation of current time-series foundation models as general-purpose time-series forecasters and highlights the need to further scale them or to fine-tune them for specific domains. It also suggests that accurately measuring the performance of foundation models can be difficult for scientific machine learning tasks, because strategies like parroting can effectively game both short- and long-term accuracy metrics.
Finding a simple but effective baseline for a challenging task can encourage rethinking of the status quo, motivating the development of better model architectures [88]. For example, context parroting formalizes an explicit baseline to compare against in the time-series domain and can help discover beyond-parroting strategies. Identifying zero-shot learning strategies beyond parroting can spur the development of next-generation foundation models and contribute to the debate on whether (or to what extent) large language models are stochastic parrots [89, 90, 91, 92].
An interesting future direction is to generalize context parroting to deal with non-stationary time series while keeping the simplicity and efficiency of the method. A promising avenue involves combining parroting with a probabilistic model, such as a Gaussian Process, to account for nonstationary trends. Additionally, the diversity of long-term forecasts can be improved by allowing stochastic selection among multiple candidate forecasts. Once generalized, the non-stationary parroting method can replace Naive and Seasonal Naive to serve as a more informative baseline for the zero-shot forecasting of general time series (weather, traffic, finance, etc.).
Finally, we want to emphasize that we are not proposing to replace time-series foundation models with context parroting. Instead, the value of parroting is as a simple baseline that can reveal the gaps in current foundation models and guide the design of future ones. When foundation models under perform relative to context parroting, it reveals that they haven’t learned to fully utilize the context data. For example, a common failure mode we observed across a range of leading foundation models (TimesFM, TimeMoE, Chronos Bolt) is that they tend to underestimate oscillations in the dynamics and the predictions often quickly converge to the mean (Fig.˜6). Being aware of context parroting as a baseline also guides the design of more informative benchmarks in the future. For example, for dynamical systems in dysts, one can generate test context data that are impossible to parrot by using only initial conditions that lead to trajectories without nearly repeating motifs.
@@ Appendix A Sample predictions from foundation models
Figure 6: Example forecasts on a chaotic system from foundation models reveal common failure modes. This is the same task as presented in Fig.˜1 (predicting the $x$ variable of the Lorenz system based on a short context trajectory with $512$ data points). Chronos does extremely well with a parroting strategy. The other models perform comparatively poorly and all exhibit a tendency to underestimate the oscillations (e.g., by quickly converging towards the mean). This is a general trend that we consistently observe across different chaotic systems and initial conditions.
@@ Appendix B Benchmarking with other metrics
Figure 7: Context parroting outperforms foundation models in zero-shot forecasting. Same setup as in Fig.˜2, but with the forecast error measured by MSE (left) and MAE (right).
Figure 8: Correlation dimension accuracy for parroting and foundation models. Each point represents the predicted correlation dimension of a chaotic attractor by a model (median of 2020 predictions from random initial conditions). The accuracy is measured by the Spearman correlation $r$ between the $135$ predicted correlation dimensions and the true correlation dimensions.
Figure 9: Valid prediction time as a function of the context length. Dashed lines indicate context lengths exceeding the maximum context window of the model. Results are averaged over $135$ chaotic systems in the dysts database, with 2020 trajectories from random initial conditions for each system.
Appendix C Effects of embedding dimension $D$
Fig.˜10 investigates how the choice of the embedding dimension $D$ affects the performance of context parroting. Overall, the valid prediction time stays consistent over a wide range of embedding dimension $D$. For short context windows, there is a slight advantage to small $D$. For long context windows, larger embedding dimensions are marginally better. This observation suggests potential improvements in the future that choose $D$ adaptively based on factors such as context length.
Figure 10: Effect of the embedding dimension $D$ on the forecast accuracy of context parroting. Each curve represents a different context length. Results are averaged over $135$ chaotic systems in the dysts database, with $100$ trajectories from random initial conditions for each system.
Appendix D Predicting scaling coefficient with the attractor dimension
Figure 11: Inverse scaling coefficient $1/\alpha$ versus correlation dimension $d_{\mathrm{cor}}$ of the chaotic attractor. We estimate $\alpha$ from the $\ell$ vs $L$ plot obtained with context parroting at $D=10$, where $L$ ranges from $2^{6}$ to $2^{16}$. This is done separately for each chaotic system. Thus, each dot represents one of the $135$ chaotic systems included in our benchmark. The Spearman rank-order correlation coefficient is $0.51\pm 0.08$ (bootstrapped standard error), and a linear regression with standard error range is underset (shaded interval). The dashed line corresponds to an exact $1:1$ scaling.
Appendix E Effects of noise and sampling rate
We add Gaussian noise to normalized chaotic trajectories and repeat the original experiments (a noise level of 0.1 translates to $10\%$ perturbation on each data point on average). The results are consistent across different orders of magnitude in noise, and parroting is consistently the best or the second best in all experiments.
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 4: Valid prediction time across noise levels (higher is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 5: MAE @ 1 Lyapunov Time across noise levels (lower is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 6: MSE @ 1 Lyapunov Time across noise levels (lower is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
$10^{-3}$ 0.73 0.85 0.52 0.58 0.52
$10^{-2}$ 0.63 0.77 0.51 0.55 0.45
$10^{-1}$ 0.59 0.57 0.37 0.49 0.16
Table 7: Correlation dimension accuracy (Spearman correlation) across noise levels (higher is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 8: KL Divergence between predicted and true attractors across noise levels (lower is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
In the main text, we set an intermediate granularity of 30 points per Lyapunov time. Below we compare it with results obtained for granularities of 10 points per Lyapunov time and 50 points per Lyapunov time. Granularity does not strongly affect the results or relative model ranking. Parroting is either the best or the second best in all experiments. This makes sense, as we would expect changing granularity to have a similar effect as rescaling of time (although with bigger or smaller gaps between data points). For example, if we use finer granularity by a factor of 2, then we would need to double the context length to get the same lookback window.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 9: Valid prediction time across different granularities (higher is better). Bold = best, italic = second best.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 10: MAE @ 1 Lyapunov Time across different granularities (lower is better). Bold = best, italic = second best.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 11: MSE @ 1 Lyapunov Time across different granularities (lower is better). Bold = best, italic = second best.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
10 0.87 0.82 0.34 0.39 0.36
30 0.80 0.85 0.52 0.58 0.52
50 0.89 0.86 0.41 0.60 0.56
Table 12: Correlation dimension accuracy (Spearman correlation) across different granularities (higher is better). Bold = best, italic = second best.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
10 0.087±0.137 0.127±0.173 0.573±0.307 0.444±0.326 0.467±0.368
30 0.122±0.194 0.173±0.209 0.346±0.297 0.345±0.298 0.354±0.290
50 0.137±0.207 0.230±0.256 0.406±0.305 0.361±0.323 0.370±0.338
Table 13: KL Divergence between predicted and true attractors across different granularities (lower is better). Bold = best, italic = second best.
Appendix F Context parroting
F.1 Discrete-token parroting